OpenStudy (anonymous):

Simplify and Write the final product in standard form and show your work 2x4(4x2 + 3x + 1)

OpenStudy (turingtest):

some rules to apply here: distribution\[a(b+c)=ab+ac\]and\[x^a\cdot x^b=x^{a+b}\]

OpenStudy (turingtest):

so for example, if we had\[3x^2(4x-7x^3+2)\]it would simplify to\[(3x^2)(4x)-(3x^2)(7x^3)+2(3x^2)\]\[(4\cdot3)(x^{2+1})-(3\cdot7)(x^{2+3})+(2\cdot3)x^2\]\[12x^3-21x^5+6x^2\]your problem works the exact same way